OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (tkhunny):

What have you tried? Give us something to go on?

OpenStudy (anonymous):

ive tried nothing

OpenStudy (tkhunny):

Well, that's no good. You MUST have learned or heard of SOME sort of technique to solve this. Throw me a bone...

OpenStudy (anonymous):

do i have to multiply?

OpenStudy (anonymous):

to start

OpenStudy (tkhunny):

Maybe. I'm tempted to multiply the second equation by -1. Try it.

OpenStudy (anonymous):

how uhh how would i set that up?

OpenStudy (tkhunny):

It's already set up. You just have to do it. x - 2y = 14 x + 3y = 9 Multiply second equation by -1 x - 2y = 14 -x - 3y = -9 That's it. You could have done that.

OpenStudy (anonymous):

36?

OpenStudy (tkhunny):

36 what? What did you do next? This is what "show your work" is all about. x - 2y = 14 -x - 3y = -9 One thing at a time. What's next? SHOW IT! Don't try to leap all the way to the end.

OpenStudy (anonymous):

i did -x - 3y = -9 and got 36

OpenStudy (tkhunny):

That makes no sense. You do not know the values for x and y. Where did the other equation go? x - 2y = 14 -x - 3y = -9 I multiplied the second equation by -1 for a reason. ADD the two equations. x - x = 0 -27 - 3y = What? 14 - 9 = What?

OpenStudy (anonymous):

im so confused i dont get this at all

OpenStudy (tkhunny):

Do you understand x - x = 0?

OpenStudy (anonymous):

yeah

OpenStudy (tkhunny):

Typo on the next one. Sorry about that. How about -2y - 3y = -5y -- Make sense?

OpenStudy (anonymous):

what do you do for the y

OpenStudy (tkhunny):

Here's where we were. x - 2y = 14 -x - 3y = -9 Think about the vertical arrangement. The x in one is over the x in the other. The y in one is over the y in the other. The constant term on the other side of "=" is lined up. Do you see these relationships?

OpenStudy (anonymous):

alright now lets do the question

OpenStudy (tkhunny):

We're trying. Do you see the relationships I have described?

OpenStudy (anonymous):

yes i think so

OpenStudy (tkhunny):

Perfect. Now, consider the two stacked equations as three separate addition problems. Add the x terms. Add the y terms. Add the constant terms. You should get three results - one for each vertical problem.

OpenStudy (anonymous):

which 3 do i do

OpenStudy (tkhunny):

The three vertical problems, as I described them in the last post.

OpenStudy (tkhunny):

Have you heard of the "Substitution Method"?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but tbh i kind of need the answer the test might time out

OpenStudy (tkhunny):

If you have not EVER heard of ANY method to solve this problem, you should get it wrong so that you can receive proper placement for your next class. It is not fair for you to have been given such a problem with NO expectation of solving it - unless this is a placement test. If this is a placement test - or any other kind of test - you should not be asking for help in this forum.